Make-or-Buy, Traditional Analysis

Morrill Company produces two different types of gauges: a density gauge and a thickness gauge. The segmented income statement for a typical quarter follows.

Density
Gauge
Thickness
Gauge

Total
Sales $ 202,500 $ 108,000 $ 310,500
Less variable expenses 108,000 62,100 170,100
  Contribution margin $ 94,500 $ 45,900 $ 140,400
Less direct fixed expenses* 27,000 51,300 78,300
Segment margin $ 67,500 $ (5,400) $ 62,100
Less common fixed expenses 40,500
Operating income $ 21,600
* Includes depreciation.

The density gauge uses a subassembly that is purchased from an external supplier for $25 per unit. Each quarter, 2,700 subassemblies are purchased. All units produced are sold, and there are no ending inventories of subassemblies. Morrill is considering making the subassembly rather than buying it. Unit-level variable manufacturing costs are as follows:

Direct materials $2
Direct labor 3
Variable overhead 2

No significant non-unit-level costs are incurred.

Morrill is considering two alternatives to supply the productive capacity for the subassembly.

  1. Lease the needed space and equipment at a cost of $36,450 per quarter for the space and $13,500 per quarter for a supervisor. There are no other fixed expenses.
  2. Drop the thickness gauge. The equipment could be adapted with virtually no cost and the existing space utilized to produce the subassembly. The direct fixed expenses, including supervision, would be $51,300, $10,800 of which is depreciation on equipment. If the thickness gauge is dropped, sales of the density gauge will not be affected.

Required:

Enter the relevant costs of each alternative.

Lease and Make Buy Drop Thickness Gauge and Make
Total relevant costs $ $ $
Answer #1

ANSWER

Total relevant costs are the incremental costs incurred

Lease and Make = (2+3+2)*2,700 + 36,450 + 13,500

Lease and Make = $68,850

Buy = 25*2,700 = $67,500

Drop Thickness Gauge and Make:

Variable cost = (2+3+2)*2,700 = $18,900

Additional Fixed Expenses = $51,300 – 10,800 = $40,500

Less: Segment loss avoided on Thickness Gauge = $5,400

Relevant cost = $54,000

Note: Depreciation is a sunk cost and is not relevant

Hence, It should make the subassembly

The alternative chosen shall be Drop Thickness Gauge and Make

------------------------------------------------------------------------

************THANK YOU******************